OpenStudy (anonymous):

Write a quadratic equation with integer coefficients and roots -2/3 and 1 ¾ A. 14x^2+13x-12=0 B.14x^2-13x-12=0 C. 12x^2-13x-14 D. 12x^2+13x-14 E.none of these. my answer is E. I am getting 6x^2-39x-7=0.Is it right?

OpenStudy (anonymous):

none of these is the right one

OpenStudy (anonymous):

(x+2/3)(x-7/4)=0 x^2 -7/4 x +2/3 x - 7/6=0 x^2 -13/12 x -7/6=0 12x^2-13x-14=0